OpenStudy (anonymous):

Find the difference quotient for f(x) = x^2 - 2x

OpenStudy (anonymous):

\(\large \frac{f(x+h)-f(x)}{h}\) since f(x)=x^2-2x, then: \(\large \frac{f(x+h)-f(x)}{h}=\frac{[(x+h)^2-(x+h)]-[x^2-x]}{h}\) can you simplify from here?

OpenStudy (anonymous):

no. what happens from there?

OpenStudy (anonymous):

simplify.... expand the binomial in the numerator and get rid of all the parenthesis and combine like terms...

OpenStudy (anonymous):

okay, thank you!

OpenStudy (anonymous):

ok.... yw...:)